# Post-detection inference for sequential changepoint localization

### v4

@@ 5 Confidence set for pre- and post-change parameters
For estimating the post-change parameter, one natural idea might be to consider the samples starting from $\max\mathcal{C}$ till the detection time $\tau$. But since both starting and ending times are random here, this strategy does not yield valid coverage guarantees. Instead, we fix each $t\in\mathcal{C}$ and construct a confidence sequence based on $X_{t},\cdots,X_{\tau}$ for $\theta_{1}$, denoted as $\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})$ (recall the definition in (4.2)), for some $\eta_{1}\in(0,1)$, whose value is less than the target non-coverage, $\alpha^{\prime}$. Finally, the union of these CS across $t\in\mathcal{C}$, yields the post-detection confidence interval for $\theta_{1}$:
$\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*}).$ (5.1)
@@
Analogously, for the pre-change parameter $\theta_{0}$, we fix each $t\in\mathcal{C}$ and construct a confidence interval for $\theta_{0}$ based on the data $X_{1},\ldots,X_{t-1}$}, denoted as $\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})$, for some $\eta_{0}\in(0,1)$, chosen to be smaller than the target (conditional) non-coverage level, $\alpha^{\prime\prime}$. The final confidence set for $\theta_{0}$ is $\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$ Next, we formalize its coverage guarantee, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods.
Theorem 5.1.
Let $\mathcal{C}$ be a confidence set, e.g., (3.9) or (4.4), such that for any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},T\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ satisfying Assumption 3.5 and $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, we have: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha$. Then, for any $T\in\mathbb{N}$,
@@ 6 Experiments
Known pre- and post-change (Setting I): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$, with $T$ being the only unknown parameter. We employ the CUSUM detector, $\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ with $A=1000$. After detection, we construct both the proposed confidence sets (universal (3.5) and adaptive (4.1)), with $B=N=100$, $\alpha=0.05$. Fig. 6.1 visualizes the confidence sets (4.1) and point estimates (3.1) across $5$ runs. Table 6.2 has the (conditional) average size and coverages of the confidence sets, the (conditional) average absolute deviation of $\hat{T}$, and delay of detection, i.e, $\tau-T$, across $500$ runs.
Figure 6.1: Setting I: The first T1T-1 observations are drawn from $N(0,1)$ and the rest from $N(1,1)$. The point estimates (3.1) are shown in a vertical red dashed line, and confidence sets (adaptive (4.1)) are shown in red points, with $B=N=100$, $\alpha=0.05,L=\infty$. Results of $5$ independent simulations are shown.
Table 6.2: Setting I: Known pre- and post-change.
@@
Universal Adaptive Universal Adaptive (Conditional) (Conditional)
100 1000 0.98 0.95 15.63 12.34 2.85 13.97
500 1000 0.98 0.95 15.77 12.57 2.62 13.22
Known pre- and parametric post-change (Setting II): We perform experiments with $F_{0}=N(0,1)$ (known pre-change), post-change class $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$, ; we vary $\Theta_{1}$ and $T$. The post-change samples are drawn from $N(1,1)$ (unknown). For detection, we use a weighted CUSUM detector:
$\tau_{\text{wcusum}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}dW(\theta)\geq A\right\}$ (6.1)
with $A=1000$, a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized mixture over $\mathcal{P}_{1}$. We use
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\theta_{0}}(X_{i})},$ (6.2)
where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$. After detection, we construct both the proposed confidence sets for $T$ (universal (3.5) with $\alpha=0.075$ and adaptive (4.3) [^5] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section C.1 (and Section D.1) of the Supplementary. with $\alpha=0.05,\beta=0.025$) with target conditional coverage $0.925$. For implementing (3.5) , we use $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})},$ where $\theta_{1}^{*}$} is the closest element of $\Theta_{0}$ from $\Theta_{1}$, and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w$ being the same weight mentioned above. In Table 6.3, we report the (conditional) average size and coverages, absolute deviation of $\hat{T}$ , and detection delay $\tau-T$ across $500$ runs. Fig. 6.4(a) shows the confidence sets (3.5) and point estimates (S2) across $5$ random runs.
Table 6.3: Known pre-change but unknown post-change. The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column. (Setting II)
@@
Universal Adaptive Universal Adaptive (Conditional) (Conditional)
100 $0.75,\infty)$ 0.98 0.97 22.21 25.72 3.95 16.87
100 $0.9,\infty)$ 0.97 0.97 17.85 24.11 3.67 16.21
500 $0.75,\infty)$ 0.98 0.97 22.89 30.83 3.45 17.63
500 $0.9,\infty)$ 0.97 0.98 18.14 28.56 3.48 15.81
@@
$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta^{*}_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (6.3)
with $A=1000$, $\theta^{*}_{0}$} is the “closest” element of $\Theta_{0}$ from $\Theta_{1}$, and for the post-change parameter, the same discrete weight distribution as in setting II is considered. We define $\hat{\theta}_{0,1:j}$ to be the MLE of $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$. We use
$\hat{T}=\operatorname*{arg\,max}_{1\leq j\leq\tau}\prod_{i=j}^{\tau}\frac{f_{\hat{\theta}_{1,j:\tau}}(X_{i})}{f_{\hat{\theta}_{0,1:j-1}}(X_{i})},$ (6.4)
where $\hat{\theta}_{1,j:n}$ is the MLE of $\theta_{1}\in\Theta_{1}$ based on $X_{j},\cdots,X_{n}$ and $\hat{\theta}_{0,:j-1}$ is the MLE of $\theta_{0}\in\Theta_{0}$ based on $X_{1},\cdots,X_{j-1}$. After detection, we construct both the proposed confidence sets (universal (3.9) with $\alpha=0.1$ and adaptive (4.4) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage $0.9$. For implementing (3.9), we use $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})}w^{\prime}(\theta)d\theta,$ with $w^{\prime}(\theta)$ being the discrete distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta-d,\theta_{3}=\theta-2d,\cdots,\theta_{10}=\theta-9d$, $d=0.2$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}w(\theta)d\theta$, with $w(\theta)$ being the same as mentioned in Setting II. $\theta_{i}^{*}$ is the closest element of $\Theta_{i}$ from the other space. Fig. 6.4(b) visualizes the confidence sets (3.9) and point estimates (S2) across 5 random runs. Table 6.5 reports the results averaged across $500$ runs.
Figure 6.4: First T1T-1 samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1),T=100$. The point estimates (S2) are shown in a vertical red dashed line. Confidence sets (universal (3.9)) are shown in red points. $N=100$, $\alpha=0.1$. Results of $5$ random simulations are shown.
Table 6.5: Confidence interval for $T$ (Setting III): unknown pre- and post-change. The true (unknown) values of the parameters are $\theta_{0}=0$, $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns.
@@
Universal Adaptive Universal Adaptive (Cond.) (Cond.)
100 $(\infty,0.25]$ $0.75,\infty)$ 0.98 0.97 26.91 26.84 4.36 25.81
100 $(\infty,0.1]$ $0.9,\infty)$ 0.97 0.97 18.63 25.71 4.19 23.15
500 $(\infty,0.25]$ $0.75,\infty)$ 0.98 0.97 26.12 31.45 4.03 24.48
500 $(\infty,0.1]$ $0.9,\infty)$ 0.97 0.98 18.86 30.66 4.07 23.07
Known pre- and nonparametric post-change (Setting IV): Here, the pre-change samples are from known $F_{0}=\text{U}[0,1]$ (uniform distribution on $[0,1]$) and let the post-change class be $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[X]\leq\mu\right\},$ where $\mathcal{M}$ is the set of all probability measures on $([0,1],\mathcal{B}([0,1]))$, with $\mu=0.25$. For detection, we use an e-detector [28] $\tau_{e}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}\geq A\right\},$ where $\hat{p}_{i-1}^{\prime}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=t}^{i-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},$ for $x\in[0,1]$, with $A=1000$. Post-detection, we construct the confidence sets (3.5). The forward $t$-delay e-process $\{R^{(t)}_{n}\}_{n\geq t}$, we use numeraire e-variable for $\mathcal{P}_{1}$ against $F_{0}$, as in Section 5.2 of [17] to get $R^{(t)}_{n}=\prod_{i=t}^{n}(1+\lambda^{*}(X_{i}-\mu)),$ where $\lambda^{*}\in(0,1/\mu)$ is the solution of the equation $e^{\lambda^{*}}=\frac{1+\lambda^{*}(1-\mu)}{1-\lambda^{*}\mu}.$ For the backward e-processes $\{S^{(t)}_{n}\}_{n<t}$, we adopt the predictable plug-in method [20]. We divide the interval $[0,1]$ into $10$ equal-sized bins $B_{i}=[\frac{i-1}{10},\frac{i}{10}),i=1,\cdots,10$ and perform histogram-based estimation. We construct $S^{(t)}_{n}=\prod_{i=n}^{t-1}\frac{\hat{p}_{i-1}(X_{i})}{{p}_{0}(X_{i})},$ where ${p}_{0}(x)=\frac{1}{10},$ and $\hat{p}_{i-1}(x)=\frac{1+\sum_{k=1}^{10}\mathds{1}(x\in B_{k})\sum_{j=i-1}^{t-1}\mathds{1}(X_{j}\in B_{k})}{9+t-i},x\in[0,1]$. We use the point estimator
$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}}\frac{\hat{p}_{i-1}^{\prime}(X_{i})}{{p}_{0}(X_{i})}.$ (6.5)
Table 6.6 shows average results across $500$ runs and Fig. 6.7(a) visualizes across $5$ runs.
Table 6.6: Setting IV: known pre-change ($\text{U}[0,1]$) but nonparametric post-change.
T
True (unknown) post-
change density, $f_{1}(x)$
Coverage
Coverage
Size
$|\hat{T}-T|$
Delay
100 $4(1-x)^{3}$ 0.98 0.97 24.14 4.06 20.65
100 $0.8\text{U}[0,0.2]+0.2\text{U}[0.2,1]$ 0.98 0.97 20.92 2.86 14.08
500 $4(1-x)^{3}$ 0.98 0.94 24.63 4.11 19.57
500 $0.8\text{U}[0,0.2]+0.2\text{U}[0.2,1]$ 0.99 0.93 21.27 2.94 15.78
@@
Composite parametric pre- and nonparametric post-change (Setting V): Here, we consider the true (unknown) pre- and post-change distributions to be $\text{N}(1,1)$ and $\text{U}[-1.2,0.8]$ respectively. Let the pre-change class be $\mathcal{P}_{0}=\left\{\text{N}(\theta,1):\theta\geq 0.5\right\},$ and the post-change class be the set of all SubGaussians with a nonpositive mean: $\mathcal{P}_{1}:=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[e^{\lambda X}]\leq e^{-\lambda^{2}/2},\forall\lambda\in[0,\infty)\right\},$ $\mathcal{M}$ is the set of all probability measures on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. We use an e-detector [28],
$\tau_{e}^{\prime}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}\geq A\right\},$
@@
with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}-0.5.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 3.5 with $P_{0}^{*}=\text{N}(0.5,1)$. After detection, we construct the confidence sets (3.9). The forward $t$-delay e-process we use is $\{R^{(t)}_{n}\}_{n\geq t}=\prod_{i=t}^{n}e^{\lambda_{i}X_{i}-\lambda_{i}^{2}/2}$, where $\lambda_{i}=\max\{0.5,\frac{1}{i-t}\sum_{j=t}^{i-1}X_{i}\}$ and the backward e-processes is constructed similarly: $\{S^{(t)}_{n}\}_{n<t}=\prod_{i=n}^{t}e^{\mu_{i}X_{i}-\mu_{i}^{2}/2}$, where $\mu_{i}=\min\{0,\frac{1}{t-i}\sum_{j=i+1}^{t}X_{j}\}-0.5$ (see, e.g., Ch. 6.7.5 of [22]). We use the following point estimator:
$\hat{T}=\operatorname*{arg\,max}_{1\leq t\leq\tau_{e}}\prod_{i=t}^{\tau_{e}^{\prime}}e^{\nu_{i}X_{i}-\nu_{i}^{2}/2}.$ (6.6)
@@
Figure 6.7: Vertical red dashed lines show the point estimates (6.6). The confidence set (3.9) is marked in red points. Results of $5$ independent simulations are shown.
Table 6.8: Setting V: Parametric pre-change, nonparametric post-change.
@@
100 0.98 0.97 27.65 3.08 27.34
500 0.98 0.94 28.03 3.15 25.12
Huber-robust (nonparametric) pre- and post-change (Setting VI):
We perform experiments for the robust Gaussian mean-change problem under the $\epsilon$ contamination model, with the true changepoint at $T=100,500$. The pre- and post-change classes are $\epsilon$ neighbourhoods around $N(\mu_{0},1)$ and $N(\mu_{1},1)$, respectively: $\mathcal{P}_{i}=\{(1-\epsilon)N(\mu_{i},1)+\epsilon F:F\in\mathcal{M}\};\penalty 10000\ \penalty 10000\ i=0,1.$ In our experiment, $\mu_{0}=0,\mu_{1}=1$ and $\epsilon=0.01,0.001$. The true (unknown) pre- and post-change distributions are $(1-\epsilon)\times N(0,1)+\epsilon\times\text{Cauchy}(-1,10)$ and $(1-\epsilon)\times N(1,1)+\epsilon\times\text{Cauchy}(-1,10)$ respectively. We use Huber’s LFD (which are also log-optimal e-values [24]) to construct a CUSUM-style change-detector as below:
$\tau=\inf\left\{t\in\mathbb{N}:\max_{1\leq j\leq t}\prod_{i=j}^{t}\pi(X_{i})\geq A\right\},$ (6.7)
where $\pi(x)=\min\{\max\{\frac{f_{1}(x)}{f_{0}(x)},c^{\prime}\},c^{\prime\prime}\}$ for constants $c^{\prime\prime}>c^{\prime}\geq 0$ as mentioned in [13, 24]. We set the threshold at $A=1000$. $\hat{T}$ is as defined in (3.3) with the same e-process, $\prod_{i=1}^{t}\pi(X_{i})$, as mentioned above. After detection, we construct our universal confidence sets (3.9) for $T$, with $R_{n}^{(t)}=1/\prod_{i=t}^{n}\pi(X_{i})$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}\pi(X_{i})$, for $n<t$ and $\alpha=0.05$. We report the average size and coverage rates across $200$ independent runs in Table 6.9. We also report the average absolute deviation of $\hat{T}$ and delay (given no false detection).
Table 6.9: Setting VI: Huber-robust (nonparametric) pre- and post-change
@@
Conditional
coverage
Marginal
coverage
Size
$|\hat{T}-T|$
Delay
100 0.01 0.98 0.98 16.74 3.23 18.38
100 0.001 0.96 0.96 15.57 3.18 13.96
500 0.01 0.97 0.93 16.18 3.87 16.58
500 0.001 0.97 0.92 15.13 3.25 14.44

### v6

@@ 5 Confidence set for pre- and post-change parameters
For estimating the post-change parameter, one natural idea might be to consider the samples starting from $\max\mathcal{C}$ till the detection time $\tau$. But since both starting and ending times are random here, this strategy does not yield valid coverage guarantees. Instead, we fix each $t\in\mathcal{C}$ and construct a confidence sequence based on $X_{t},\cdots,X_{\tau}$ for $\theta_{1}$, denoted as $\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*})$ (recall the definition in (18)), for some $\eta_{1}\in(0,1)$, whose value is less than the target non-coverage, $\alpha^{\prime}$. Finally, the union of these CS across $t\in\mathcal{C}$, yields the post-detection confidence interval for $\theta_{1}$:
$\cup_{t\in\mathcal{C}}\operatorname{CS}(X_{t},\cdots,X_{\tau};1-\eta_{1}r_{t}^{*}).$ (21)
@@
Analogously, for the pre-change parameter $\theta_{0}$, we fix each $t\in\mathcal{C}$ and construct a confidence interval for $\theta_{0}$ based on the data $X_{1},\ldots,X_{t-1}$}, denoted as $\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*})$, for some $\eta_{0}\in(0,1)$, chosen to be smaller than the target (conditional) non-coverage level, $\alpha^{\prime\prime}$. The final confidence set for $\theta_{0}$ is $\cup_{t\in\mathcal{C}}\operatorname{CI}(X_{1},\cdots,X_{t-1};1-\eta_{0}r_{t}^{*}).$
Next, we formalize its coverage guarantee, which is nonasymptotic and agnostic to the detection algorithm, unlike existing methods.
Theorem 9.
Let $\mathcal{C}$ be a confidence set, e.g., (14) or (20), such that for any $\theta_{0}\in\Theta_{0},\theta_{1}\in\Theta_{1},T\in\mathbb{N}$, change detection algorithm $\mathcal{A}$ satisfying Assumption 1 and $\mathbb{\mathbb{P}}_{\theta_{0},\infty}(\tau\geq T)\neq 0$, we have: $\mathbb{P}_{\theta_{0},T,\theta_{1}}(T\in\mathcal{C}\mid\tau\geq T)\geq 1-\alpha$. Then, for any $T\in\mathbb{N}$,
@@ 6 Experiments
We validate our methods across various simulation and real-data settings, demonstrating their practicality.
Known pre- and post-change (Setting I): Here, we draw $X_{1},\cdots,X_{T-1}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,1)$ and $X_{T},X_{T+1},\cdots\stackrel{{\scriptstyle iid}}{{\sim}}N(1,1)$, with $T$ being the only unknown parameter. We employ the CUSUM detector, $\tau=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\prod_{i=j}^{n}\frac{f_{1}(X_{i})}{f_{0}(X_{i})}\geq A\right\},$ with $A=1000$. After detection, we construct both the proposed confidence sets (universal (10) and adaptive (17)), with $B=100,N=50$, $\alpha=0.1$. Fig. 1 visualizes the confidence sets (17) across $5$ runs. Table 2 has the (conditional on the event $\tau\geq T$) average size and coverages of the confidence sets and delay of detection, i.e, $\tau-T$, across $500$ runs.
Figure 1: Setting I. The first T1T-1 observations are drawn from $N(0,1)$ and the rest from $N(1,1)$. The confidence sets (adaptive (17)) are shown in red points, with $B=N=100$, $\alpha=0.1,L=\infty$. Results of 55 independent simulations are shown.
Table 2: Setting I: Known pre- and post-change.
@@
Universal Adaptive Universal Adaptive (Conditional)
100 1000 0.948 0.906 12.412 9.694 12.803
500 1000 0.942 0.908 12.176 9.278 13.227
Known pre- and parametric post-change (Setting II): We perform experiments with $F_{0}=N(0,1)$ (known pre-change), post-change class $\mathcal{P}_{1}=\{N(\theta,1):\theta\in\Theta_{1}\}$, ; we vary $\Theta_{1}$ and $T$. The post-change samples are drawn from $N(1,1)$ (unknown). For detection, we use the CUSUM detector with $A=1000$, a discrete weight distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta+d,\theta_{3}=\theta+2d,\cdots,\theta_{10}=\theta+9d$, $d=0.2$, $\theta=0.75$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$, which appeared to us to be a sensible discretized mixture over $\mathcal{P}_{1}$. After detection, we construct both the proposed confidence sets for $T$ (universal (10) with $\alpha=0.075$ and adaptive (19) [^4] Implementation details of the adaptive method in Setting II (and Setting III) are the same as discussed in Section C.1 (and Section D.1) of the Supplementary. with $\alpha=0.1,\beta=0.025$) with target conditional coverage $0.925$. For implementing (10) , we use $R_{n}^{(t)}=\prod_{i=t}^{n}\frac{f_{\theta_{0}}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})},$ where $\theta_{1}^{*}$} is the closest element of $\Theta_{0}$ from $\Theta_{1}$, and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}}(X_{i})}w(\theta)d\theta$, with $w$ being the same weight mentioned above. In Table 3, we report the results averaged across $500$ runs. Fig. 2(a) shows the confidence sets (10) across $5$ random runs.
Table 3: Known pre-change but unknown post-change. The pre-change parameter $\theta_{0}=0$ is known, the true (unknown) post-change parameter is $\theta_{1}=1$, while the algorithm only knows $\theta_{1}\in\Theta_{1}$, which we mention in the second column. (Setting II)
@@
Universal Adaptive Universal Adaptive (Conditional)
100 $0.75,\infty)$ 0.972 0.966 16.870 14.824 13.878
100 $0.9,\infty)$ 0.976 0.972 12.860 11.552 12.212
500 $0.75,\infty)$ 0.969 0.968 16.168 14.266 13.638
500 $0.9,\infty)$ 0.972 0.971 14.734 12.602 12.814
@@
$\tau_{\text{wcs-ripr}}=\inf\left\{n\in\mathbb{N}:\max_{1\leq j\leq n}\int_{\theta\in\Theta_{1}}\prod_{i=j}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta^{*}_{0}}(X_{i})}dW(\theta)\geq A\right\},$ (22)
with $A=1000$, $\theta^{*}_{0}$} is the “closest” element of $\Theta_{0}$ from $\Theta_{1}$, and for the post-change parameter, the same discrete weight distribution as in setting II is considered. We define $\hat{\theta}_{0,1:j}$ to be the MLE of $\theta_{0}\in\Theta_{0}$ based on data $X_{1},\cdots,X_{j}$. After detection, we construct both the proposed confidence sets (universal (14) with $\alpha=0.1$ and adaptive (20) with $\alpha=0.05,\beta=\gamma=0.025$) for $T$, with $N=B=100$, both having target conditional coverage $0.9$. For implementing (14), we use $R_{n}^{(t)}=\int_{\theta\in\Theta_{0}}\prod_{i=t}^{n}\frac{f_{\theta}(X_{i})}{f_{\theta_{1}^{*}}(X_{i})}w^{\prime}(\theta)d\theta,$ with $w^{\prime}(\theta)$ being the discrete distribution taking values $\theta_{1}=\theta,\theta_{2}=\theta-d,\theta_{3}=\theta-2d,\cdots,\theta_{10}=\theta-9d$, $d=0.2$, when $\mathcal{P}_{1}=\{N(\mu,1):\mu>\theta\}$, with exponentially decaying weights $w_{i}=e^{-\frac{i-1}{2}}-e^{-\frac{i}{2}},\text{ for }i=1,\cdots,9,\text{ and }w_{10}=e^{-9/2}$ and $S_{n}^{(t)}=\int_{\theta\in\Theta_{1}}\prod_{i=n}^{t-1}\frac{f_{\theta}(X_{i})}{f_{\theta_{0}^{*}}(X_{i})}w(\theta)d\theta$, with $w(\theta)$ being the same as mentioned in Setting II. $\theta_{i}^{*}$ is the closest element of $\Theta_{i}$ from the other space. Fig. 2(b) visualizes the confidence sets (14) across 5 random runs. Table 4 reports the results averaged across $500$ runs.
Figure 2: First T1T-1 samples are drawn from $N(0,1)$ and the remaining samples from $N(1,1),T=100$. Confidence sets (universal (14)) are shown in red points. $N=100$, $\alpha=0.1$. Results of 55 random simulations are shown.
Table 4: Confidence interval for $T$ (Setting III): unknown pre- and post-change. The true (unknown) values of the parameters are $\theta_{0}=0$, $\theta_{1}=1$, while the algorithm only knows $\theta_{i}\in\Theta_{i},i=0,1$, which are mentioned in the second and third columns.
@@
Universal Adaptive Universal Adaptive (Conditional)
100 $(\infty,0.25]$ $0.75,\infty)$ 0.984 0.982 23.683 23.827 22.810
100 $(\infty,0.1]$ $0.9,\infty)$ 0.968 0.978 15.552 15.294 15.282
500 $(\infty,0.25]$ $0.75,\infty)$ 0.982 0.976 23.825 24.482 24.060
500 $(\infty,0.1]$ $0.9,\infty)$ 0.970 0.942 16.860 16.956 15.576
@@
Composite parametric pre- and nonparametric post-change (Setting IV): Here, we consider the true (unknown) pre- and post-change distributions to be $\text{N}(1,1)$ and $\text{U}[-1.2,0.8]$ respectively. Let the pre-change class be $\mathcal{P}_{0}=\left\{\text{N}(\theta,1):\theta\geq\mu_{1}\right\},\mu_{1}=0.75$ and the post-change class be the set of all SubGaussians with a nonpositive mean: $\mathcal{P}_{1}=\left\{F\in\mathcal{M}:\mathbb{E}_{F}[e^{\lambda X}]\leq e^{-\frac{\lambda^{2}}{2}},\forall\lambda\in[0,\infty)\right\}$, $\mathcal{M}$ is the set of all probability measures on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. We use an e-detector [shin2022detectors],
$\tau_{e}^{\prime}=\inf\left\{n\in\mathbb{N}:\max_{1\leq t\leq n}\prod_{i=t}^{n}e^{(\nu_{i}-\mu_{1})X_{i}-\nu_{i}^{2}/2+\mu_{1}^{2}/2}\geq A\right\},$
@@
with $A=1000$, $\nu_{i}=\min\{0,\frac{1}{i-1}\sum_{j=1}^{i-1}X_{j}\}.$ This $\mathcal{P}_{0}$ and $\tau_{e}^{\prime}$ satisfies Assumption 1 with $P_{0}^{*}=\text{N}(\mu_{1},1)$. After detection, we construct the confidence sets (14). The forward $t$-delay e-process we use is $\{R^{(t)}_{n}\}_{n\geq t}=\prod_{i=t}^{n}e^{\lambda_{i}X_{i}-\lambda_{i}^{2}/2}$, where $\lambda_{i}=\max\{\mu_{1},\frac{1}{i-t}\sum_{j=t}^{i-1}X_{i}\}$ and the backward e-processes is constructed similarly: $\{S^{(t)}_{n}\}_{n<t}=\prod_{i=n}^{t}e^{(\eta_{i}-\mu_{1})X_{i}-\eta_{i}^{2}/2+\mu_{1}^{2}/2}$, where $\eta_{i}=\min\{0,\frac{1}{t-i}\sum_{j=i+1}^{t}X_{j}\}$ (see, e.g., Ch. 6.7.5 of [ramdas2024hypothesis]). Table 5 contains the results across $500$ runs and Fig. 3 visualizes across $5$ runs.
@@
Figure 3: The confidence set (14) is marked in red points. Results of 55 independent simulations are shown.
Table 5: Setting IV: Parametric pre-change, nonparametric post-change.
@@
100 0.984 0.982 23.859 9.959
500 0.992 0.930 25.381 9.606
Huber-robust (nonparametric) pre- and post-change (Setting V):
We perform experiments for the robust Gaussian mean-change problem under the $\epsilon$ contamination model, with the true changepoint at $T=100,500$. The pre- and post-change classes are $\epsilon$ neighbourhoods around $N(\mu_{0},1)$ and $N(\mu_{1},1)$, respectively: $\mathcal{P}_{i}=\{(1-\epsilon)N(\mu_{i},1)+\epsilon F:F\in\mathcal{M}\};\penalty 10000\ \penalty 10000\ i=0,1.$ In our experiment, $\mu_{0}=0,\mu_{1}=1$ and $\epsilon=0.01,0.001$. The true (unknown) pre- and post-change distributions are $(1-\epsilon)\times N(0,1)+\epsilon\times\text{Cauchy}(-1,10)$ and $(1-\epsilon)\times N(1,1)+\epsilon\times\text{Cauchy}(-1,10)$ respectively. We use Huber’s LFD (which are also log-optimal e-values [saha2024huber]) to construct a CUSUM-style change-detector as below:
$\tau=\inf\left\{t\in\mathbb{N}:\max_{1\leq j\leq t}\prod_{i=j}^{t}\pi(X_{i})\geq A\right\},$ (23)
where $\pi(x)=\min\{\max\{\frac{f_{1}(x)}{f_{0}(x)},c^{\prime}\},c^{\prime\prime}\}$ for constants $c^{\prime\prime}>c^{\prime}\geq 0$ as mentioned in [huber1965robust, saha2024huber]. We set the threshold at $A=1000$. After detection, we construct our universal confidence sets (14) for $T$, with $R_{n}^{(t)}=1/\prod_{i=t}^{n}\pi(X_{i})$, for $n\geq t$ $S_{n}^{(t)}=\prod_{i=n}^{t-1}\pi(X_{i})$, for $n<t$ and $\alpha=0.1$. We report the results averaged across $500$ independent runs in Table 6, and Fig. 3 visualizes the results across $5$ runs.
Table 6: Setting V: Huber-robust (nonparametric) pre- and post-change
@@
Conditional coverage
Conditional size
Conditional delay
100 0.01 0.939 13.298 14.140
500 0.01 0.948 13.634 13.943
100 0.001 0.944 12.274 12.638
500 0.001 0.958 11.927 12.496